# User Experience in Dataset Search Platform Interfaces

### v1

@@ 4.2.6. Feedback and Help
P29: 68% participants indicated that they could easily locate the help documentation, and the information related to the functionality was easily discovered and understood. N63: 47% participants observed that while they can easily find general information like ”Contact Us” and ”Documentation”, they are often not interested in other options such as ”Publications” and ”Events” present in the navigation bar when searching for datasets. Their focus is primarily on dataset-related features.
4.3. New Prototypes for Dataset Search Platform
4.3.1. Initial Interaction
@@
Figure 3. First Page of A Dataset Search Tool
Prototype of First Page of Dataset Search Tool
Figure 3. First Page of A Dataset Search Tool
4.3.2. Search Process
@@
To enhance user interaction and platform efficiency, several improvements are recommended. Immediate auto-suggestions and spelling corrections could be provided through real-time prompts like ”did you mean.” Dataset search platform could also incorporate Natural Language Processing techniques such as entity extraction and synonym recognition for more accurate results. Special attention should be paid to varying user needs; for instance, new users might appreciate a help query option in the search bar, while other users may prioritize datasets based on factors like geographic location (Shown in Figure  4). Additionally, implementing common search operators like ”AND,” ”OR,” and ”NOT” can refine search specificity. To make these operators more accessible, they could be visualized in the search bar interface, accompanied by dropdown menus or clickable icons. An ’add more’ option could also be provided for scalable keyword and operator input. This approach not only informs users about advanced search tools but also enables them to generate highly targeted results from the beginning (Shown in Figure  5).
Figure 4. Search Bar Considering Varied User Needs
Prototype of Search Bar Considering Varied User Needs
Figure 4. Search Bar Considering Varied User Needs
Figure 5. Search Bar Considering Enhanced Precision
Prototype of Search Bar Considering Enhanced Precision
Figure 5. Search Bar Considering Enhanced Precision
4.3.3. Dataset Exploration
@@
Dataset previews need to be brief but informative. They should underscore search terms, date of publication, quality of metadata, view counts, licence type, file dimensions, covered regions, and the nature of the dataset. Additionally, licence information should be streamlined to quickly signify if commercial use is permitted (See Figure  6). The previews should also be succinct while offering depth, including expandable segments for extra details. These segments can feature key terms, descriptions of columns, and ratios of missing values, presented in a bullet-point format (See Figure  7).
Figure 6. User-Focused Dataset Previews Section
Prototype of Dataset Preview Considering Prioritizing User Preferences
Figure 6. User-Focused Dataset Previews Section
Figure 7. Expandable Section For Additional Information
Expandable Section For Additional Information about dataset metadata
Figure 7. Expandable Section For Additional Information
Figure 8. Simplified Result Display Section
Prototype of Optimizing Result Previews for Enhanced User Experience
Figure 8. Simplified Result Display Section
@@
Given the numerous dataset previews, users can face reading fatigue. To counteract this, a colour-coded scheme could be employed to highlight datasets meeting user-defined criteria. For example, using dark green to signify criteria that match the user’s definitions, intermediate blue for criteria that don’t match, and pale orang for criteria that the user hasn’t defined (Shown in Figure  8), while being accessible to colour-blind users, tested in Section  4.3.5.
4.3.4. Usage of Filters and Sorting
@@
To enhance the user experience across platforms, several improvements are suggested. Filters should allow both single and multiple selections, symbolised by circles and boxes respectively. These filters should encompass a range of options including time range, dataset size, and license type. Ideally positioned at the top of the page, filters should not necessitate a full page refresh upon each selection. Instead, a subtle bottom-right corner notification should indicate the number of datasets that fit the selected criteria (Shown in Figure  9).
Figure 9. Improved Filter Design
Prototype of Filters Considering Optimizing Placement and Interactivity
Figure 9. Improved Filter Design
@@
Sorting should employ dual criteria: initial sorting based on relevance, followed by secondary sorting according to publication date. These improvements aim for a more efficient and user-friendly dataset search process, with subsequent figures outlining the proposed changes (Shown in Figure  10).
Figure 10. Dual Criteria Sorting
Prototype of Dual Criteria Sorting
Figure 10. Dual Criteria Sorting
4.3.5. Dataset Actions
@@
Improvements in dataset actions could be multifaceted. First, the citation functionalities across platforms should be expanded to include diverse styles like LaTeX, catering to academic requirements. Second, implementing a feature to allow users to directly compare datasets could be invaluable. The attributes of datasets should be colour-coded for more straightforward comparisons. A dark green fill could signify attributes where another dataset performs better than the currently selected one, a intermediate blue fill for areas where the current dataset excels, and a pale orange fill to highlight unique features not offered by the selected dataset. This visual guide would streamline the decision-making process, rendering it more intuitive and less cognitively taxing for users (Shown in Figure  11).
Figure 11. Datasets Comparison Page
Prototype of Datasets Comparison Page
Figure 11. Datasets Comparison Page
@@
The researcher used the Coblis Colour Blindness Simulator (col, 2023) to test 8 types of colour blindness on the Figure  11 provided . The results show clear variations in colour fills for each type of colour blindness, ensuring accessibility for those with colour vision deficiencies (Shown in Figure  12).
Figure 12. Eight Colour Blindness Tests
Coblis Colour Blindness Simulator
Figure 12. Eight Colour Blindness Tests
4.3.6. Feedback and Help
@@
For enhancements, both platforms could incorporate introductory tutorials to familiarize new users with their respective features. A step-by-step guide, necessitating users to acknowledge each tutorial segment before advancing, can ensure a more comprehensive understanding of the platform’s functionalities. By instituting this level of guidance, platforms can increase user engagement, reduce initial learning curves, and subsequently, elevate the overall user experience (Shown in Figure  13).
Figure 13. Tutorial Illustration
Prototype of Tutorial Illustration
Figure 13. Tutorial Illustration
5. Discussion and Limitations
@@
6. Conclusion and Future Works
@@
Acknowledgements.
The author wishes to acknowledge the contributions of supervisors Professor Elena Simperl and Dr. Albert Meroño Peñuela for their insightful guidance and inspiration. Thanks are also directed to Bohui Zhang for editorial assistance. Gratitude is extended to the varied participants of the user study and the Department of Informatics at King’s College London for providing essential research materials and support in participant recruitment.

### v2

@@ 4.1 Google Dataset Search
Lastly, in feedback and help, 68% of participants found the help documentation easy to locate and understand (P29). However, 47% noted that while general information like "Contact Us" and "Documentation" was easy to find, options such as "Publications" and "Events" were often irrelevant when focused on searching for datasets (N63).
4.3 New Prototypes for Dataset Search Platform
4.3.1 Initial Interaction
@@
Figure 3: First Page of A Dataset Search Tool
4.3.2 Search Process
@@
To enhance user interaction and platform efficiency, several improvements are recommended. Immediate auto-suggestions and spelling corrections could be provided through real-time prompts like "did you mean." Dataset search platform could also incorporate Natural Language Processing techniques such as entity extraction and synonym recognition for more accurate results. Special attention should be paid to varying user needs; for instance, new users might appreciate a help query option in the search bar, while other users may prioritize datasets based on factors like geographic location (Shown in Figure  4). Additionally, implementing common search operators like "AND," "OR," and "NOT" can refine search specificity. To make these operators more accessible, they could be visualized in the search bar interface, accompanied by dropdown menus or clickable icons. An ’add more’ option could also be provided for scalable keyword and operator input. This approach not only informs users about advanced search tools but also enables them to generate highly targeted results from the beginning (Shown in Figure  5).
Figure 4: Search Bar Considering Varied User Needs
Figure 5: Search Bar Considering Enhanced Precision
4.3.3 Dataset Exploration
@@
Dataset previews need to be brief but informative. They should underscore search terms, date of publication, quality of metadata, view counts, licence type, file dimensions, covered regions, and the nature of the dataset. Additionally, licence information should be streamlined to quickly signify if commercial use is permitted (See Figure  6). The previews should also be succinct while offering depth, including expandable segments for extra details. These segments can feature key terms, descriptions of columns, and ratios of missing values, presented in a bullet-point format.
Figure 6: User-Focused Dataset Previews Section
Figure 7: Simplified Result Display Section
@@
Given the numerous dataset previews, users can face reading fatigue. To counteract this, a colour-coded scheme could be employed to highlight datasets meeting user-defined criteria. For example, using dark green to signify criteria that match the user’s definitions, intermediate blue for criteria that don’t match, and pale orang for criteria that the user hasn’t defined (Shown in Figure  7), while being accessible to colour-blind users, tested in Section  4.3.5.
4.3.4 Usage of Filters and Sorting
@@
To enhance the user experience across platforms, several improvements are suggested. Filters should allow both single and multiple selections, symbolised by circles and boxes respectively. These filters should encompass a range of options including time range, dataset size, and license type. Ideally positioned at the top of the page, filters should not necessitate a full page refresh upon each selection. Instead, a subtle bottom-right corner notification should indicate the number of datasets that fit the selected criteria (Shown in Figure  8).
Figure 8: Improved Filter Design
@@
Sorting should employ dual criteria: initial sorting based on relevance, followed by secondary sorting according to publication date. These improvements aim for a more efficient and user-friendly dataset search process.
4.3.5 Dataset Actions
@@
Improvements in dataset actions could be multifaceted. First, the citation functionalities across platforms should be expanded to include diverse styles like LaTeX, catering to academic requirements. Second, implementing a feature to allow users to directly compare datasets could be invaluable. The attributes of datasets should be colour-coded for more straightforward comparisons. A dark green fill could signify attributes where another dataset performs better than the currently selected one, a intermediate blue fill for areas where the current dataset excels, and a pale orange fill to highlight unique features not offered by the selected dataset. This visual guide would streamline the decision-making process, rendering it more intuitive and less cognitively taxing for users (Shown in Figure  9).
Figure 9: Datasets Comparison Page
@@
The researcher used the Coblis Colour Blindness Simulator [1] to test 8 types of colour blindness on the Figure  9 provided . The results show clear variations in colour fills for each type of colour blindness, ensuring accessibility for those with colour vision deficiencies (Shown in Figure  10).
Figure 10: Eight Colour Blindness Tests
4.3.6 Feedback and Help
@@
For enhancements, both platforms could incorporate introductory tutorials to familiarize new users with their respective features. A step-by-step guide, necessitating users to acknowledge each tutorial segment before advancing, can ensure a more comprehensive understanding of the platform’s functionalities. By instituting this level of guidance, platforms can increase user engagement, reduce initial learning curves, and subsequently, elevate the overall user experience (Shown in Figure  11).
Figure 11: Tutorial Illustration
5 Discussion and Limitations
@@
6 Conclusion and Future Works